The refrigerant R-134a (CF3CH2F, tetrafluorethane) has a specific volume υs=0.0309 m3/kg at a pressure p=0.26 MPa.

B. Calculate the molar mass of tetrafluorethane CF3CH2F.

MR-134a = ????? g

C. Determine the temperature T of tetrafluorethane.

T =   ???? K
Answer #1

Answer:

B) molar mass M is a physical property defined as the mass of a given substance from the chemical elements.

SI unit for molar mass is kg/mol.

CF3CH2F: F atomic weight = 19

C atomic weight = 12

H atomic weight =1

CF3CH2F: =12 + (3 x19) + 12 + (2x1) + 19

= 12 + 57 + 12 + 2 + 19

= 102

C) P = 0.26 Mpa = 2.566 atm (1Mpa = 9.869 atm)

Specific volume = 0.0309 m3/ kg

PV =nRT

PV = (wt/mwt) RT wt = weight

mwt = molecular weight (CF3CH2F mwt = 102)

P(V/wt)= RT/mwt V/wt = 0.0309 m3/ kg (according to the question)

R = gass constant (8.314 joules per Kelvin (K) per mole)

2.566 x 0.0309 = 8.314 x T / 102

T = 2.566 x 0.0390 x 102/ 8.314

T = 1.227 K